OpenStudy (anonymous):

Very simple question, extremely simple like basic algebra. What would I do to find x in this e^x+1=10

OpenStudy (anonymous):

e^x = 9 x*ln(e) = ln 9 x*1 = ln 9 x = ln 9

OpenStudy (anonymous):

ln e is always 1